40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue May 15 22:11:38 2018, gautam, Update, General, IFO maintenance OL_RIN_2018_05_15.pdfOLsums.png
    Reply  Fri Jun 15 09:14:42 2018, Steve, Update, General, Oplev sums opSums.png
Message ID: 13971     Entry time: Fri Jun 15 09:14:42 2018     In reply to: 13847
Author: Steve 
Type: Update 
Category: General 
Subject: Oplev sums 

Oplev sums of 240 days.

Quote:

Since there have been various software/hardware activity going on (stack weighing, AUX laser PLL, computing timing errors etc etc), I decided to do a check on the state of the IFO.

  • c1susaux, c1aux and c1iscaux crates were keyed as they were un-telnet-able.
  • Single arm locking worked fine, TT alignment was tweaked (as these had drifted due to the ADC failure in c1lsc) to maximize Y arm transmission using the dither servos.
  • Arms weren't staying locked for extended periods of time. I particularly suspected ITMX, as I saw what I judged to be excess motion on the Oplev.
  • @Steve - ITMX and BS HeNes look like they are in need of replacement judging by the RIN (although the trend data doesn't show any precipitous drop in power). If we are replacing the BS/PRM Oplev HeNe, might be a good time to plan the inejction path a bit better on that table.
  • RIN in Attachment #1 has been normalized by the mean value of the OL sum channel. There is now a script to make this kind of plot from NDS in the scripts directory (as I found it confusing to apply different calibrations to individual traces in DTT).

 

Attachment 1: opSums.png  36 kB  Uploaded Fri Jun 15 10:16:10 2018  | Hide | Hide all
opSums.png
ELOG V3.1.3-